OpenStudy (kambol1898):

solve for x x-70/x<-3

RhondaSommer (rhondasommer):

Could you screenshot the question? Id be happy to help :D \[\Large\color{cyan}{\mathbb{WELCOME\\TO\\OPENSTUDY}}\]

OpenStudy (misty1212):

\[\frac{x-70}{x}<-3\] or \[x-\frac{70}{x}<-3\]?

RhondaSommer (rhondasommer):

Thats why I wanted a screenshot ^_^

RhondaSommer (rhondasommer):

about* you ARE on a study site after all

OpenStudy (kambol1898):

@misty1212 its the second equation